1. Speed is 72 km/h. Convert km/h to m/s.

Ans.:



1
Expert's answer
2021-09-27T09:01:27-0400

"1km = 1000m"

"1h = 3600s"

"72\frac {km}{h}= \frac{72*1000}{3600}\frac{m}{s}= 20\frac{m}{s}"

"\text{Answer: }20\frac{m}{s}"